# Fast Real Evaluation Through Sound Mixed-Precision Tuning

### v1

@@ 4. Fast Sound Precision Tuning
Reval’s precision assignment algorithm is based on error bounds for floating-point computations. At a high level, these error bounds are based on error Taylor series and condition numbers, with some variations to ensure soundness and performance. Their derivation is fairly technical; we invite readers uninterested in the technical details to focus on the high-level description in Section 4.1 and then on Equations (1), (2), and (3), which are the key results proved in the rest of this section.
4.1. High-level Approach
Intuitively, the endpoints computed in interval arithmetic bound all possible faithfully-rounded evaluations of a given expression. Sound bounds on faithfully-rounded evaluations thus indicate how wide that interval may be. If that error bound is parameterized by the precision of each operation, then the sound error bound is an inequality that we can solve for an appropriate precision assignment. Importantly, these error bounds need to be sound, always over-approximating the error; otherwise, a too-low precision will be assigned and the resulting output interval will be too wide. But if the error is over-approximated by too much, too-high a precision will be assigned, which will make evaluation take too long.
The need for soundness and accuracy suggests error Taylor series, but performance is a problem. Error Taylor series are only sound if the “higher-order error” term is included, and that term is expensive to compute: it requires computing quadratically-many second derivatives of the expression. We instead introduce a novel variation on error Taylor series that achieves soundness using only first derivatives. Equation 1 shows that variation. Even this variation, computed naively, would take $O(n^{2})$ time to compute (where $n$ is the number of real operations), but an algebraic reorganization inspired by (and related to) condition numbers allows computing it in $O(n)$ time. That reorganization is shown in Equation 2. Finally, the error bound can be solved for an appropriate precision assignment; the solution is shown in Equation 3.
@@ 4.2. Sound First-Order Error
Analogously to error Taylor series, consider a fixed sequence $\hat{z}_{i}=\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$, of floating-point operations $\hat{f}_{i}$ on input and output floating-point registers $\hat{x}_{i}$, $\hat{y}_{i}$, and $\hat{z}_{i}$; we assume each $\hat{x}_{i}$ is itself the output register $\hat{z}_{j}$ of some prior operation. Naturally, some operations have one or even zero arguments (for floating-point constants like $\pi$), but for simplicity we write $\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$ for each operation. Each operation also has a corresponding real-number operation $z_{i}=f(x_{i},y_{i})$. We also assume each $\hat{f}_{i}$ is faithfully rounded,888 Reval’s interval arithmetic is not tight, especially in challenging cases like $\cos(\pi)$, but adding a couple of extra bits allows for a couple of ULPs of looseness; see Section 6 for more. so that
@@
for some $|\varepsilon_{i}|\leq 2^{-p_{i}}$, where $p_{i}$ is the precision used for this operation.999 This formula is not correct in the presence of under- and over-flow, though many cases of such are handled using movability flags (Flatt and Panchekha, 2023), a feature of the interval library we use. The remaining cases are handled by the slack mechanism of Section 6. Substituting in $\hat{x}_{i}$ and $\hat{y}_{i}$, which are themselves computed in prior instructions, yields $\hat{z}_{i}$ as a function of all of the $\varepsilon_{j}$s, which we now write $\hat{z}_{i}(\vec{\varepsilon})$.
The relative error of $\hat{z}_{i}$, then, is a function of the $\vec{\varepsilon}$. Error Taylor series provide a way to bound this relative error. Notice that $\hat{z}_{i}(\vec{0})=z_{i}$; the numerator of the relative error formula is thus $\hat{z}_{i}(\vec{\varepsilon})-\hat{z}_{i}(\vec{0})$. But $\hat{z}_{i}(\vec{0})$ is the first term of the Taylor series for $\hat{z}_{i}$.
Here we diverge from error Taylor series, applying the Lagrange remainder theorem at order 1 instead of order 2: [^10] There are many variations of the Lagrange remainder theorem; the statement below is one of them but can also be seen as a corollary of the mean value theorem.
@@
for some $0\leq\vec{\varepsilon^{*}}\leq\vec{\varepsilon}\leq 2^{-p_{i}}$ (or the symmetrical bound for $\varepsilon_{j}<0$). In this bound, $\vec{\varepsilon}$ represents the rounding error of a faithfully-rounded execution, so $\vec{\varepsilon^{*}}$ represents the rounding error of some other, also faithfully-rounded execution (of $\partial f/\partial\varepsilon_{j}$).
We don’t know the signs of the $\varepsilon_{i}$, just a bound on their magnitude, but using the triangle inequality we can still bound the relative error of $\hat{z}_{i}$:
@@ 4.3. Linear-Time Error Bounds
$+(1+2^{-p_{i}})\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\left|\frac{y_{i}(\partial_{2}f_{i})(x^{*}_{i},y^{*}_{i})}{f_{i}(x_{i},y_{i})}\right|\right)}_{\operatorname{ampl}_{2}(\hat{z}_{i})}\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\sum_{j<i}2^{-p_{j}}\left|\frac{1}{y_{i}}\frac{\partial\hat{y_{i}}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|\right)}_{\operatorname{bound}(\hat{y}_{i})}$
@@
The intuition behind this formula is that every operation $\hat{z}_{i}$ introduces some error $\operatorname{intro}(\hat{z}_{i})$ and also amplifies error in its inputs by factors of $\operatorname{ampl}_{k}(\hat{z}_{i})$. The error bound for $\hat{z}_{i}$ just sums these three factors. Readers familiar with condition numbers will immediately see the similarity between Equation 2 and the condition number formula $e_{z}=\mu+\Gamma_{x}e_{x}+\Gamma_{y}e_{y}$ (Zou et al., 2019). However, condition numbers bound only the first-order error so are unsound, while Equation 2 is sound. This in turn causes differences like the $\operatorname{intro}(\hat{z}_{i})$ term in place of $\mu$, the slight difference in definition between $\operatorname{ampl}_{k}(\hat{z}_{i})$ and $\Gamma_{k}$, and the extra $1+2^{-p_{i}}$ term. These differences are small but necessary for soundness. [^11] In fact, the name $\operatorname{ampl}$ is a pun on both “amplification factor” and ”ample” as in sound.
@@ 5. The Exponent Trick
To actually use Equation 2 we must compute $\operatorname{intro}(\hat{z}_{i})$ and $\operatorname{ampl}_{k}(\hat{z}_{i})$, which requires optimizing over $\varepsilon^{*}$. However, we can use interval arithmetic to perform this optimization. Moreover, already-computed interval outputs from each operation can be reused to make this fast.
@@ 5.1. Optimizing Over $\vec{\varepsilon^{*}}$
$\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{f(x_{i}^{*},y_{i}^{*})}{f(x_{i},y_{i})}\right|\right\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{z_{i}^{*}}{z_{i}}\right|\right\rceil\in\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil$
@@
The upper bound on $\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil$ is the high endpoint of $\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil$. Write $\operatorname{lo}(\bar{x})$ and $\operatorname{hi}(\bar{x})$ for the high and low endpoint of an interval $\bar{x}$, and assume for now (see Section 6) that $\bar{x}$ contains neither zero nor infinity. Then,
@@
We call this the “exponent trick”: bounding the logarithms of $\operatorname{intro}$ and $\operatorname{ampl}_{k}$ using only: 1) integer operations, 2) on $\operatorname{minlog}$ and $\operatorname{maxlog}$, 3) of already-computed intervals $\bar{x}$, $\bar{y}$, and $\bar{z}$. The remainder of this section describes exponent tricks for various operators. Some of the derivations are mathematically challenging, but all the key results are shown in Tables 2 and 2.
@@ 5.2. Simple amplification factors
Division, square and cube roots, exponents, and logarithms allow a similar algebraic rearrangement, producing a formula that uses only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}_{i}$, $\bar{y}_{i}$, and $\bar{z}_{i}$. The resulting formulas are shown in Table 2.
@@ 5.3. Approximations
However, for some operators, an approximation has to be performed to bound $(\partial_{k}f)(x,y)$ via some expression over $x$, $y$, or $z$. This approximation allows us to construct an exponent tricks formula for $\operatorname{ampl}_{k}(f)$, that is, a formula that bounds $\lceil\log_{2}\operatorname{ampl}_{k}(f)\rceil$ using only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}$, $\bar{y}$, and $\bar{z}$. Table 2 shows all of the approximations used and the exponent trick formulas they result in, and readers are invited to focus their attention on it. The rest of this section merely contains detailed derivations for each row of that table.
@@
$\lceil\log_{2}\operatorname{ampl}_{1}(\sin(\hat{x}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{y}.$
@@
$\lceil\log_{2}\operatorname{ampl}_{1}(\cos(\hat{z}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{y}+\min(\operatorname{maxlog}(\bar{x}),0).$
These approximations are sound and easy to compute. Moreover, the approximation error is small when either the input or the output is near zero, which helps reduce the needed precision.
@@
The power function $x^{y}$ has an amplification factor for $y$ that uses a $\log(x^{*})$ term. This means that we must bound $\log_{2}|\log(x)|$ in terms of $\log_{2}|x|$. To do so, we first rewrite
@@
The amplification factors for $\sinh$ and $\cosh$ both require bounding $\tanh(x)$. In both cases, we use the well-known fact that $|\tanh(x)|\leq\min(|x|,1)$, leading to $\lceil\log_{2}|\tanh(x)|\rceil\leq\min(\operatorname{maxlog}(\bar{x}),0)$. This approximation is relatively tight, producing at most one bit of error.
The amplification factor for $\tanh(x)$ requires bounding $x/\sinh(x)\cosh(x)$. We use a very weak bound, $x/\sinh(x)\cosh(x)\leq 1$. This over-approximates significantly for large values of $x$, though not near 0.
@@
which now only uses already-computed intervals. Moreover, this approximation is off by at most one bit.
@@
Operator ($f$) Approximation used & exponent tricks bound
@@
$\operatorname{maxlog}(\bar{x})+\operatorname{maxlog}(\bar{y})-2\max(\operatorname{minlog}(\bar{x}),\operatorname{minlog}(\bar{y}))-\operatorname{minlog}(\bar{z})$
@@
6. Implementation Details
Reval provides a simple user interface: the user first compiles an expression, which may have free variables, and then applies it to specific values. The compilation produces a simple register machine with a vector of instructions and separate vectors with, for each instruction, an output interval and a precision assignment.
@@ 6.1. Compilation
Reval supports expressions with real-number operations, and real constants, boolean operations, and if statements, but no loops. Specifically, Reval supports the constants $\pi$ and $e$, the standard arithmetic operators (plus, minus, times, divide), and every function specified in the POSIX standard for math.h, including square and cube roots; exponents, powers, and logarithms; trigonometric, arc-trigonometric, and hyperbolic and arc-hyperbolic functions; the error and gamma functions; modulus and remainder operators; rounding functions; and miscellaneous others like $\min$ and $\max$. It also supports comparison functions and boolean connectives. Thanks to the lack of loops, each expression can be flattened to a linear instruction sequence. Users can also supply a batch of multiple expressions to evaluate on the same inputs. Reval performs common subexpression elimination so this can dramatically reduce runtime if the expressions have shared structure, though this feature is not used in this paper.
Instructions are written in two-argument form, reading in their arguments from input registers and returning a result that is saved in the instruction’s output register. Interval operators come from the Rival library (Flatt and Panchekha, 2023), which has a couple of unusual features. It provides boolean functions like comparisons and connectives that output ternary “boolean intervals”; for example, $[1,1]<[0,2]$ returns the “maybe” interval [,]bottomtop[$\bot$,$\top$][ ⊥ , ⊤ ]. An if statement with a “maybe” condition unions the intervals in both branches. In the presence of domain errors, Rival intervals enclose all outputs on valid inputs; for example, $\sqrt{[-1,1]}=[0,1]$. A special empty interval $\emptyset$ is used when the whole input interval is invalid. All intervals also track domain errors using a ternary error flag; for $\sqrt{[-1,-1]}$ the error flag is $\top$ and for $\sqrt{[-1,1]}$ the error flag is [,]bottomtop[$\bot$,$\top$][ ⊥ , ⊤ ]. Rival also provides movability flags (Flatt and Panchekha, 2023) that detect some unevaluable inputs.
Reval typically attempts to correctly round each expression to double precision. However, this rounding behavior is customizable by supplying a discretization. A discretization is a type $T$, a target precision $t$, a conversion function from MPFR values to $T$s, and a distance function on pairs of $T$s. The distance function is called with the two converted $T$ endpoints of an interval and returns one of three values: the endpoints can be equal, unequal, or neighbors, meaning that the two $T$ values share a rounding boundary. The “equal” return value is used for correct rounding; if the user instead wants faithful rounding, they can supply a distance function that returns “neighbor” instead of “equal”.

### v2

@@ 4. Fast Sound Precision Tuning
Reval’s precision assignment algorithm is based on error bounds for floating-point computations. Intuitively, the endpoints computed in interval arithmetic bound all possible faithfully-rounded evaluations of a given expression. Sound bounds on faithfully-rounded evaluations thus indicate how wide that interval may be. If that error bound is parameterized by the precision of each operation, then the sound error bound is an inequality that we can solve for an appropriate precision assignment. Ideally the bounds are not only sound but also relatively tight, to avoid over-estimating the necessary precision.
The need for soundness and accuracy suggests error Taylor series, but performance is a problem. Error Taylor series are only sound if the “higher-order error” term is included, which involves the quadratically-many second derivatives of the expression. We introduce a slight variation on error Taylor series that achieves soundness using only first derivatives; Equation 1 shows that variation. An algebraic reorganization inspired by (and related to) condition numbers allows computing this variation in $O(n)$ time; Equation 2 shows that reorganization. Finally, the error bound can be solved for an appropriate precision assignment; Equation 3 shows that solution. The derivations of each equation are fairly technical and we invite readers to skip them and jump directly equations on their first read-through.
@@ 4.1. Sound First-Order Error
Analogously to error Taylor series, consider a fixed sequence $\hat{z}_{i}=\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$, of floating-point operations $\hat{f}_{i}$ on input and output floating-point registers $\hat{x}_{i}$, $\hat{y}_{i}$, and $\hat{z}_{i}$; we assume each $\hat{x}_{i}$ is itself the output register $\hat{z}_{j}$ of some prior operation. Naturally, some operations have one or even zero arguments (for floating-point constants like $\pi$), but for simplicity we write $\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$ for each operation. Each operation also has a corresponding real-number operation $z_{i}=f(x_{i},y_{i})$. We also assume each $\hat{f}_{i}$ is faithfully rounded,777 Reval’s interval arithmetic is not tight, especially in challenging cases like $\cos(\pi)$, but adding a couple of extra bits allows for a couple of ULPs of looseness; see Section 6 for more. so that
@@
for some $|\varepsilon_{i}|\leq 2^{-p_{i}}$, where $p_{i}$ is the precision used for this operation.888 As before, under- and over-flow are handled using movability flags (Flatt and Panchekha, 2023), a feature of the interval library we use. The remaining cases are handled by the slack mechanism of Section 6. Substituting in $\hat{x}_{i}$ and $\hat{y}_{i}$, which are themselves computed in prior instructions, yields $\hat{z}_{i}$ as a function of all of the $\varepsilon_{j}$s, which we now write $\hat{z}_{i}(\vec{\varepsilon})$.
The relative error of $\hat{z}_{i}$, then, is a function of the $\vec{\varepsilon}$. Error Taylor series provide a way to bound this relative error. Notice that $\hat{z}_{i}(\vec{0})=z_{i}$; the numerator of the relative error formula is thus $\hat{z}_{i}(\vec{\varepsilon})-\hat{z}_{i}(\vec{0})$. But $\hat{z}_{i}(\vec{0})$ is the first term of the Taylor series for $\hat{z}_{i}$, so we are looking to bound the error of a Taylor series approximation. To do so, we apply the Lagrange remainder theorem at order 1: [^9] There are many variations of the Lagrange remainder theorem; the statement below is one of them but can also be seen as a corollary of the mean value theorem.
@@
for some $0\leq\vec{\varepsilon^{*}}\leq\vec{\varepsilon}\leq 2^{-p_{i}}$ (or the symmetrical bound for $\varepsilon_{j}<0$). In this bound, $\vec{\varepsilon}$ represents the rounding error of a faithfully-rounded execution, so $\vec{\varepsilon^{*}}$ represents the rounding error of some other, also faithfully-rounded execution.
Since we don’t know the signs of $\hat{\varepsilon}$, we bound the relative error of $\hat{z}_{i}$ using the triangle inequality:
@@ 4.2. Linear-Time Error Bounds
$+(1+2^{-p_{i}})\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\left|\frac{y_{i}(\partial_{2}f_{i})(x^{*}_{i},y^{*}_{i})}{f_{i}(x_{i},y_{i})}\right|\right)}_{\operatorname{ampl}_{k}(\hat{z}_{i})}\underbrace{\left(\max_{\vec{\varepsilon^{*}}}\sum_{j<i}2^{-p_{j}}\left|\frac{1}{y_{i}}\frac{\partial\hat{y_{i}}}{\partial\varepsilon_{j}}(\vec{\varepsilon^{*}})\right|\right)}_{\operatorname{bound}(\hat{y}_{i})}$
@@
The intuition behind this formula is that every operation $\hat{z}_{i}$ introduces some error $\operatorname{intro}(\hat{z}_{i})$ and also amplifies error in its inputs by factors of $\operatorname{ampl}_{k}(\hat{z}_{i})$. The error bound for $\hat{z}_{i}$ just sums these three factors. Readers familiar with condition numbers will immediately see the similarity between Equation 2 and the condition number formula $e_{z}=\mu+\Gamma_{x}e_{x}+\Gamma_{y}e_{y}$ (Zou et al., 2019). However, condition numbers bound only the first-order error so are unsound, while Equation 2 is sound. This in turn causes small, but necessary, [^10] In fact, the name $\operatorname{ampl}$ is a pun on both “amplification factor” and ”ample” as in sound. differences like the $\operatorname{intro}(\hat{z}_{i})$ term in place of $\mu$, the slight difference in definition between $\operatorname{ampl}_{k}(\hat{z}_{i})$ and $\Gamma_{k}$, and the extra $1+2^{-p_{i}}$ term.
@@ 5. The Exponent Trick
To actually use Equation 2 we must compute $\operatorname{intro}(\hat{z}_{i})$ and $\operatorname{ampl}_{k}(\hat{z}_{i})$, which requires optimizing over $\varepsilon^{*}$. We can use interval arithmetic to perform this optimization. Moreover, already-computed interval outputs from each operation can be reused to make this fast.
@@ 5.1. Optimizing Over $\vec{\varepsilon^{*}}$
$\lceil\log_{2}\operatorname{intro}(\hat{z}_{i})\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{f(x_{i}^{*},y_{i}^{*})}{f(x_{i},y_{i})}\right|\right\rceil=\max_{\vec{\varepsilon^{*}}}\left\lceil\log_{2}\left|\frac{z_{i}^{*}}{z_{i}}\right|\right\rceil\in\lceil\log_{2}|\bar{z}_{i}/\bar{z}_{i}|\rceil.$
@@
Write $\operatorname{lo}(\bar{x})$ and $\operatorname{hi}(\bar{x})$ for the high and low endpoint of an interval $\bar{x}$, and assume for now (see Section 6) that $\bar{x}$ contains neither zero nor infinity. Then,
@@
We call this the “exponent trick”: bounding the logarithms of $\operatorname{intro}$ and $\operatorname{ampl}_{k}$ using only: 1) integer operations, 2) on $\operatorname{minlog}$ and $\operatorname{maxlog}$, 3) of already-computed intervals $\bar{x}$, $\bar{y}$, and $\bar{z}$. The remainder of this section describes exponent tricks for various operators’ $\operatorname{ampl}_{k}$ terms. Some of the derivations are mathematically challenging, but all the key results are shown in Tables 2 and 2.
@@ 5.2. Simple amplification factors
Division, square and cube roots, exponents, and logarithms allow a similar algebraic rearrangement, producing a formula that uses only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}_{i}$, $\bar{y}_{i}$, and $\bar{z}_{i}$. The resulting “exponent tricks” formulas are shown in Table 2.
@@ 5.3. Approximations
However, for some operators, an approximation has to be performed to bound $(\partial_{k}f)(x,y)$ via some expressions over $x$, $y$, or $z$. These approximations allow us to construct exponent tricks formulas for $\operatorname{ampl}_{k}(f)$, that is, formulas that bound $\lceil\log_{2}\operatorname{ampl}_{k}(f)\rceil$ using only integer operations on $\operatorname{maxlog}$ and $\operatorname{minlog}$ of $\bar{x}$, $\bar{y}$, and $\bar{z}$.
Table 2 shows all of the approximations used and the exponent trick formulas they result in, and readers are invited to focus their attention on it.
The rest of this section merely contains detailed derivations for each row of that table.
@@
$\lceil\log_{2}\operatorname{ampl}_{1}(\sin(\hat{x}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{z}.$
@@
$\lceil\log_{2}\operatorname{ampl}_{1}(\cos(\hat{x}))\rceil\leq\operatorname{maxlog}\bar{x}-\operatorname{minlog}\bar{z}+\min(\operatorname{maxlog}(\bar{x}),0).$
@@
The $\operatorname{ampl}_{2}$ factor for the power function $x^{y}$ has uses a $\log(x^{*})$ term. This means that we must bound $\log_{2}|\log(x)|$ in terms of $\log_{2}|x|$. To do so, we first rewrite
@@
The amplification factors for $\sinh$ and $\cosh$ both use $\tanh(x)$. Because $|\tanh(x)|\leq\min(|x|,1)$, we have $\lceil\log_{2}|\tanh(x)|\rceil\leq\min(\operatorname{maxlog}(\bar{x}),0)$. This approximation is relatively tight, producing at most one bit of error.
The amplification factor for $\tanh(x)$ requires bounding $x/\sinh(x)\cosh(x)$. We use a very weak bound, $x/\sinh(x)\cosh(x)\leq 1$, which is only tight near 0.
@@
which now only uses already-computed intervals and is at most one bit from being tight.
@@
Operator ($f$) Approximation used & exponent tricks bound for $\lceil\log_{2}\operatorname{ampl}_{k}\rceil$
@@
$\operatorname{maxlog}(\bar{x})+\operatorname{maxlog}(\bar{y})-2\min(\operatorname{minlog}(\bar{x}),\operatorname{minlog}(\bar{y}))-\operatorname{minlog}(\bar{z})$
@@
6. Implementing Real Evaluation
Reval provides a simple user interface: the user first compiles (one or more) expressions, which may have free variables, and then applies it to specific values for those free variables. The compilation produces a simple register machine with a vectors of instructions, their precision assignments, and their outputs.
@@ 6.1. Compilation
Reval supports expressions containing the constants $\pi$ and $e$, the standard arithmetic operators (plus, minus, times, divide), and every function specified in the POSIX standard for math.h, including square and cube roots; exponents, powers, and logarithms; trigonometric, arc-trigonometric, and hyperbolic and arc-hyperbolic functions; the error and gamma functions; modulus and remainder operators; rounding functions; and miscellaneous others like $\min$ and $\max$. It also supports comparison functions, boolean connectives, and if statements. Interval operators are provided by the Rival library (Flatt and Panchekha, 2023), which also provides error flags for detecting domain errors and movability flags for detecting unevaluable expressions. Reval compiles the expression(s) to a register machine instruction sequence in three-argument form and performs common subexpression elimination.
By default, Reval evaluates each expression to correctly-rounded double precision values. This rounding behavior is customizable by supplying a discretization: a type $T$, a target precision $t$, a conversion from MPFR to $T$s, and a distance function on pairs of $T$s. The distance function is called with the two converted $T$ endpoints of an interval and decides whether the endpoints are equal, unequal, or neighbors, where the last is used for correct rounding.